Car a is traveling east at 20m/s and car b is traveling east at 30m/s, if car b is 200m behind car a how long will it take for car b to catch car a?


1
Expert's answer
2021-01-10T18:28:54-0500

For moving in one direction:

"t=\frac{s}{v_2-v_1}=\frac{200}{30-20}=20~s."